OpenStudy (djbreezy):

What is the solution to the system? Line 1 x y –1 9 3 –3 Line 2 x y –1 –1 2 5 ( , )

OpenStudy (amilapsn):

just derive the equations of the lines and solve for x and y...

OpenStudy (anonymous):

(5,11)

OpenStudy (anonymous):

Point slope form again -3-9/3+1= -3 y+1=-3(x-9) y+1=-3x+27 y=-3x+26 5+1/2+1=2 y+1=2(x+1) y+1=2x+2 y=2x+1 For your system of equations you have y={y=-3x+26} y={y=2x+1} There is a couple ways you can solve this. You could either use a subtraction method, or a substitution method. since they are both y = something i would just substitute one y in one or the other equation. 2x+1=-3x+26 5x+1=26 5x=25 x=5 So now you just plug in x = 5 for each quotation and you have your y's for both solution points y=-3(5)+26 y=11 y=2(5)+1 y=11 So this works out nice, your solutions for the set of equations is (5,11) since they are the same points